OpenStudy (anonymous):

x=(2^n + 3^n + 5^n + 7^n) / 11^n n is a natural number........ Find all the possible integer value of x

OpenStudy (anonymous):

@experimentX @mukushla @Traxter

OpenStudy (anonymous):

for n>1\[2^n + 3^n + 5^n + 7^n< 11^n \]and for n=1 it becomes\[\frac{17}{11}\]so there is no integer value in the form\[\frac{2^n + 3^n + 5^n + 7^n}{ 11^n} \]

OpenStudy (anonymous):

Well try this: (2^n +3^n +5^n + 7^n +13^n)/11^n